OpenStudy (kimberly101):

Use equations to solve the following questions At a certain store, a football costs $15 more than a basketball. If 3 basketballs and 2 footballs cost a total of $115, find the cost of each ball.

OpenStudy (anonymous):

*Let: x = cost of basketball y = cost of football; *since a football costs 15 $ more, you have the first equation: y = x + 15 $ *next, if you bought 3 basketballs, and 2 footballs which costs 115 $, you will have the second equation: 115 $ = 3x + 2y *substitute the first equation to the second: 115 $ = 3x + 2(x + 15 $) 115 $ = 3x + 2x + 30$ (put together similar terms and simplify, note the change of signs once you move an expression to the other side) 115 $ - 30 $ = 3x + 2x 85 $ = 5x (divide both sides by 5 to get the value of x) 85$/5 = 5x/5 17$=x so, a basketball costs 17$. *to get the value of y/ or cost of football, substitute value of x to first equation. y = x + 15 $ y = (17$) + 15 $ y = 32 $ so, a football costs 32$. *to check the answer, substitute both values to second equation. 115 $ = 3x + 2y 115 $ = 3(17$) + 2(32$) 115 $ = 51$+ 64$ 115 $ = 115 $ that means the answer is right. got from yahoo answer

OpenStudy (kimberly101):

thankyou!!!